Find the equation of the line, in slope-intercept form, that satisfies the given conditions.
The graph is parallel to the graph of
x + 2y = 8
and passes through the point whose coordinates are
(−2, −4).

Answer by Cromlix(4381)

Hi there,
First sort equation x + 2y = 8
into y = mx + c form.
2y = -x + 8
y = -1/2x + 4.
Gradient = -1/2
Lines that are parallel to
one another are equal.
m1 = m2
Setting up new line equation.
Using formula:
y - b = m(x - a)
m = -1/2 and coords (-2, -4)
y -(-4)= -1/2(x -(-2))
y + 4 = -1/2(x + 4)
y + 4 = -1/2x - 2
y = -1/2x - 2 - 4
y = -1/2x - 6
or multiply through by 2
2y = -x - 12
or rearranging
x + 2y = -12
Hope this helps :-)